Express −1 + i in polar form. Hence show that (−1 + i)16 is real and that 1/(−1 + i)6
is purely imaginary, giving the value for each.
1
Expert's answer
2020-05-25T20:31:07-0400

"z=-1+i=|z|(cos\theta+isin\theta)"

"|z|=\sqrt{1+1}=\sqrt{2}"

"cos\theta=-1/\sqrt{2}, sin\theta=1/\sqrt{2}, \theta=3\pi/4"

"-1+i=\sqrt{2}(cos\frac {3\pi}{4}+isin\frac {3\pi}{4})"


"z^n=|z|^n(cos(n\theta)+isin(n\theta))"

"(-1+i)^{16}=(\sqrt{2})^{16}(cos(\frac {16\cdot3\pi}{4})+isin(\frac {16\cdot3\pi}{4})="

"=2^8(cos12\pi+isin12\pi)=256(1+0)=256"


"\frac {1}{(-1+i)^6}=\frac {1}{(\sqrt{2})^6(cos(\frac {6\cdot3\pi}{4})+isin(\frac {6\cdot3\pi}{4})}="

"=\frac {1}{2^3(cos(9\pi/2)+isin(9\pi/2))}=\frac {1}{8(0+i)}=\frac {1}{8i}=-i/8"